OpenStudy (anonymous):

Write as ordered pair??? y = -3x2 + 6x

OpenStudy (anonymous):

What is X?

OpenStudy (anonymous):

I don't know, I have to solve for x, I think to find ordered pair.

OpenStudy (anonymous):

Well I got 0, tbh. (−3)(x)(2)+6x =(−3)(x)(2)+6x =0

OpenStudy (anonymous):

So Y = 0

OpenStudy (anonymous):

Thanks. Saphira

OpenStudy (ybarrap):

In set notation, $$ \{(x,y):y=-3x^2+6x\} $$

OpenStudy (ybarrap):

Y does not equal 0. http://www.wolframalpha.com/input/?i=y+%3D+-3x2+%2B+6x

OpenStudy (anonymous):

ok. So how would I get the ordered pair, I still don't understand..

OpenStudy (anonymous):

Yeah how would it @ybarrap ? If y is not 0

OpenStudy (anonymous):

I feel more confused...

OpenStudy (ybarrap):

An ordered pair is a set consisting of "pairs of numbers." Ordered pairs are written like (0,1), (2,3),(-5,2), where the 1st coordinate is x and the 2nd coordinate is y. So to sample the equation you have, y=-3x^2 +6, set x=0 (for example) and see what y is: -3(0)^2+6=0+6 = 6. So one ordered pair for your equation is (0,6). This is a "Point" on the line y=-3x^2+6. To find ALL ordered pairs, you would need to find for ALL values of x the corresponding values of y. There would be an infinite number of them. The way to write this is using set-builder notation, which I did above. What that set describes is the set of all x and y such that the "points" (x,y) satisfy the equation y=-3x^2 +6. This is the easiest way to describe ALL ordered pairs.

OpenStudy (anonymous):

Ok. Thank you for your time.

OpenStudy (anonymous):

Oh x was 0 not y. I was confused

OpenStudy (anonymous):

I thought it was y lol

OpenStudy (anonymous):

no worries Saphira, you did way better than I! <3

OpenStudy (anonymous):

lol